OpenStudy (anonymous):

Determine which of the following ordered pairs is the solution of the given system. x + 3y = 10 2x - y = 6 A. (4, 2) B. (7, 1) C. (16, -2) D. (3, 0)

OpenStudy (nurali):

x + 3y = 10 2x - y = 6

OpenStudy (nurali):

x + 3y = 10---(1) 2x - y = 6 ---(2) in equation (1) x=10-3y

OpenStudy (nurali):

put x in equation (2) 2x-y=6 2(10-3y)-y=6 20-6y-y=6 20-7y=6 7y=14 y=2

OpenStudy (nurali):

put y in equation (1) x+3y=10 x+3(2)=10 x+6=10 x=4 so solution is (x,y)=(4,2)